# Compositional Difference-in-Differences for categorical outcomes

### v1

@@ A.4. Proof of Theorem 2
Thus, bounding $\Delta_{1}^{k}$ yields bounds on $q_{1,1}^{N}(c_{k})$.
Assumption 3 states:
$\Delta_{1}^{k}\in\operatorname{Conv}\left\{\Delta_{0}^{k},\Delta_{-1}^{k}\right\}=\left[\min\{\Delta_{0}^{k},\Delta_{-1}^{k}\},\;\max\{\Delta_{0}^{k},\Delta_{-1}^{k}\}\right].$
Therefore, the sharp bounds on $\Delta_{1}^{k}$ are:
$\underline{\Delta}_{1}^{k}=\min\{\Delta_{0}^{k},\Delta_{-1}^{k}\},\quad\overline{\Delta}_{1}^{k}=\max\{\Delta_{0}^{k},\Delta_{-1}^{k}\}.$
Plugging into (1), we obtain sharp bounds on $q_{1,1}^{N}(c_{k})$:
$b_{\min}^{k}:=\exp\left(\underline{\Delta}_{1}^{k}+\log(q_{0,1}^{N}(c_{k}))\right)=\exp\left(\min\left\{\log\left(\frac{q_{1,0}^{N}(c_{k})}{q_{0,0}^{N}(c_{k})}\right),\log\left(\frac{q_{1,-1}^{N}(c_{k})}{q_{0,-1}^{N}(c_{k})}\right)\right\}+\log(q_{0,1}^{N}(c_{k}))\right),$
$b_{\max}^{k}:=\exp\left(\overline{\Delta}_{1}^{k}+\log(q_{0,1}^{N}(c_{k}))\right)=\exp\left(\max\left\{\log\left(\frac{q_{1,0}^{N}(c_{k})}{q_{0,0}^{N}(c_{k})}\right),\log\left(\frac{q_{1,-1}^{N}(c_{k})}{q_{0,-1}^{N}(c_{k})}\right)\right\}+\log(q_{0,1}^{N}(c_{k}))\right).$
Assumption 4 states:
$\Delta_{1}^{k}\in\operatorname{Conv}\left\{\min_{t<0}\Delta_{t}^{k},\;\max_{t<0}\Delta_{t}^{k}\right\}=\left[\underline{\Delta}^{k},\;\overline{\Delta}^{k}\right],$
where
$\underline{\Delta}^{k}:=\min_{t<0}\Delta_{t}^{k},\quad\overline{\Delta}^{k}:=\max_{t<0}\Delta_{t}^{k}.$
Thus, the bounds are:
$b_{\min}^{k}=\exp\left(\underline{\Delta}^{k}+\log(q_{0,1}^{N}(c_{k}))\right)=\exp\left(\min_{t<0}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right)+\log(q_{0,1}^{N}(c_{k}))\right),$
$b_{\max}^{k}=\exp\left(\overline{\Delta}^{k}+\log(q_{0,1}^{N}(c_{k}))\right)=\exp\left(\max_{t<0}\log\left(\frac{q_{1,t}^{N}(c_{k})}{q_{0,t}^{N}(c_{k})}\right)+\log(q_{0,1}^{N}(c_{k}))\right).$
Assumption 5 states:
@@
Let $q_{1,1}^{I}(c_{k})$ denote the observed outcome for group 1 in period 1 (i.e., the treated outcome). The causal effect is often defined as the ratio of observed to counterfactual (in multiplicative models):
$\text{GTT}(c_{k}):=\frac{q_{1,1}^{I}(c_{k})}{q_{1,1}^{N}(c_{k})}.$
Since $q_{1,1}^{N}(c_{k})\in[b_{\min}^{k},b_{\max}^{k}]$, and the function $x\mapsto 1/x$ is decreasing for $x>0$, we have:
@@
This can be expressed via the log-ratio transformation $\ell$. Then the set of possible log-ratios satisfies:
$\log(b_{\min}^{k})-\log(b_{\max}^{p})\leq\ell(\pi)_{k}\leq\log(b_{\max}^{k})-\log(b_{\min}^{p}),$
@@
= The key idea is that the uncertainty in each component $q_{1,1}^{N}(c_{k})\in[b_{\min}^{k},b_{\max}^{k}]$ translates into a set-identified region for the share vector $\pi_{1,1}^{N}$. The bounds $b_{\min}^{k},b_{\max}^{k}]$ are sharp under each relaxation, because the convex hull assumption allows $\Delta_{1}^{k}$ to attain any value in the interval, including the endpoints. The treatment effect bounds follow directly from monotonicity of the ratio function.
A.5. Proof of Theorem 3
@@ Appendix B Adding covariates
This section extends our main identification results to incorporate covariates. Let $X\in\mathcal{X}$ denote a vector of covariates, which may be endogenous. For this analysis, we restrict our attention to the case where $X$ is discrete. In this setting, we can redefine all relevant quantities by stratifying on $X$. This yields a set of new estimands, each conditional on a specific covariate profile $x$. Provided that the untreated potential quantities $q_{g,t}^{N}(x)$ are strictly positive for all groups, time periods, and values $x\in\mathcal{X}$, we can formulate the following conditional assumption:
@@
The main challenge arises when $X$ includes continuous variables, as it becomes necessary to model or approximate the conditional distribution in a flexible way, rather than relying on simple stratification. We do not explicitly discuss this case here.
@@ Appendix C Implementation of CoDiD and Inference procedure
$\hat{g}_{1,1}^{N}=\exp\cdot\left(\log\cdot\left(\hat{q}_{1,0}^{N}\right)+\log\cdot\left(\hat{q}_{0,1}^{N}\right)-\log\cdot\left(\hat{q}_{0,0}^{N}\right)\right)$
@@
For inference, we use the parametric bootstrap to approximate the sampling distribution of the parameters. The parametric bootstrap is a resampling technique designed for situations where the data-generating process can be modeled parametrically. Instead of drawing resamples directly from the observed data, the procedure assumes a parametric model for the data and uses estimated parameters to generate new synthetic samples. For a detailed treatment of the method, see Horowitz (2019). In our case, the procedure proceeds as follows. For each bootstrap iteration $b=1,\dots,B$, where $B$ is the total number of iterations, we:
@@
    $q_{g,t}^{(b)}(c_{1},\dots,c_{p})\sim\text{Multinomial}\left(S_{g,t},\hat{\pi}_{g,t}(c_{1}),\dots,\hat{\pi}_{g,t}(c_{p})\right).$
@@
Appendix D Random utility model and log odds
Consider a standard Random Utility Model (RUM), where an individual chooses among $q$ discrete alternatives $\{c_{1},c_{2},...,c_{p}\}$. The utility associated with alternative $c_{k}$ is
$U_{k}=V_{k}+\varepsilon_{k},$
where $V_{k}$ is the deterministic component of utility (observable to the researcher), and $\varepsilon_{k}$ is the stochastic component (unobserved preference heterogeneity). The individual chooses the alternative with the highest utility:
$Choice=\arg\max_{k}U_{k}.$
The probability that alternative $k$ is chosen is
$\pi(c_{k})=P(Choice=c_{k})=\Pr(U_{j}\geq U_{k},\ \forall k\neq j),$
which can be rewritten as
$\pi(c_{k})=\Pr(\varepsilon_{k}-\varepsilon_{j}\leq V_{j}-V_{k},\ \forall k\neq j).$
Assuming the stochastic components $\varepsilon_{k}$ are independent and identically distributed as Type I Extreme Value (Gumbel), the choice probabilities take the closed form of the multinomial logit:
$\pi(c_{k})=\frac{\exp(V_{k})}{\sum_{g=1}^{p}\exp(V_{j})}.$
For any two alternatives $c_{k}$ and $c_{p}$, the odds ratio of choosing $c_{k}$ over $c_{p}$ is
$\frac{\pi(c_{k})}{\pi(c_{p}))}=\frac{\exp(V_{k})}{\exp(V_{p})}=\exp(V_{k}-V_{p}),$
and taking logarithms gives
$\log\left(\frac{\pi(c_{k})}{\pi(c_{p})}\right)=V_{k}-V_{p}$
This shows that the log-odds of choosing one alternative over another are equal to the difference in deterministic utilities, an increase in $V_{k}$ relative to $V_{p}$ increases the probability of choosing $k$.

### v2

@@ A.5. Proof of Theorem 2
Thus, bounding $\Delta_{1}^{k}$ yields bounds on $q_{1,1}^{N}(c_{k})$. Assumption 3 states:
@@
Let $q_{1,1}^{I}(c_{k})$ denote the observed outcome for group 1 in period 1 (i.e., the treated outcome). Since $q_{1,1}^{N}(c_{k})\in[b_{\min}^{k},b_{\max}^{k}]$, and the function $x\mapsto 1/x$ is decreasing for $x>0$, we have:
@@
A.6. Proof of Theorem 3
@@ Appendix B Adding covariates
This section extends the main identification results to incorporate covariates. Let $X\in\mathcal{X}$ denote a vector of covariates, which may be endogenous. For this analysis, I restrict my attention to the case where $X$ is discrete. In this setting, I can redefine all relevant quantities by stratifying on $X$. This yields a set of new estimands, each conditional on a specific covariate profile $x$. Provided that the untreated potential quantities $q_{g,t}^{N}(x)$ are strictly positive for all groups, time periods, and values $x\in\mathcal{X}$, I can formulate the following conditional assumption:
@@
The main challenge arises when $X$ includes continuous variables, as it becomes necessary to model or approximate the conditional distribution in a flexible way, rather than relying on simple stratification. I do not explicitly discuss this case here.
@@ Appendix C Implementation of CoDiD and Inference procedure
$\hat{q}_{1,1}^{N}=\exp\cdot\left(\log\cdot\left(\hat{q}_{1,0}^{N}\right)+\log\cdot\left(\hat{q}_{0,1}^{N}\right)-\log\cdot\left(\hat{q}_{0,0}^{N}\right)\right)$
@@
For inference, I use the parametric bootstrap to approximate the sampling distribution of the parameters. The parametric bootstrap is a resampling technique designed for situations where the data-generating process can be modeled parametrically. Instead of drawing resamples directly from the observed data, the procedure assumes a parametric model for the data and uses estimated parameters to generate new synthetic samples. For a detailed treatment of the method, see Horowitz (2019). In my case, the procedure proceeds as follows. For each bootstrap iteration $b=1,\dots,B$, where $B$ is the total number of iterations, we have:
@@
    $q_{g,t}^{(b)}(c_{1},\dots,c_{p})\sim\text{Multinomial}\left(S_{gt},\hat{\pi}_{g,t}(c_{1}),\dots,\hat{\pi}_{g,t}(c_{p})\right).$
@@
Appendix D General random utility model without IIA
But now, assume that the error terms jointly follow a Generalized Extreme Value (GEV) distribution,
$F(\varepsilon_{1},\dots,\varepsilon_{p})=\exp\left[-G\left(e^{-\varepsilon_{1}},\dots,e^{-\varepsilon_{p}}\right)\right],$
where $G(\cdot)$ is a homogeneous generating function that determines the correlation structure among the unobserved utility shocks. When $G(y_{1},\dots,y_{p})=\sum_{k=1}^{p}y_{k}$, the shocks are independent Gumbel, and the model collapses to the multinomial logit case. More general choices of $G$ (such as those defining the nested or cross-nested logit models) allow correlation across alternatives, thereby relaxing the independence of irrelevant alternatives (IIA) assumption while preserving closed-form choice probabilities. Train (2009) gives us different possible examples for $G$ as shown in table 6. Consider the same model (modified to account for correlation between error terms) :
$U^{N}_{gt}(c_{k})=V^{N}_{gt}(c_{k})+\varepsilon_{gt}(c_{k})=\mu^{N}_{gt}(c_{k})+\log\left(G_{k}(\mu^{N}_{gt})\right)-\log\left(G(\mu^{N}_{gt})\right)+\log\left(S^{N}_{gt}\right)+\varepsilon_{gt}(c_{k})$
Table 6. Examples of GEV Models
Model $G(\cdot)$ $\pi^{N}(c_{k})$
Logit $\sum_{j=1}^{p}e^{V^{N}(c_{j})}$ $\frac{e^{V^{N}(c_{k})}}{\sum_{j}e^{V^{N}(c_{j})}}$
Nested Logit $\sum_{m=1}^{M}\Big(\sum_{j\in\mathcal{B}_{m}}e^{V^{N}(c_{j})/\lambda_{m}}\Big)^{\lambda_{m}}$ $\frac{e^{V^{N}(c_{k})/\lambda_{m}}\Big(\sum_{j\in\mathcal{B}_{m}}e^{V^{N}(c_{j})/\lambda_{m}}\Big)^{\lambda_{m}-1}}{\sum_{r}\Big(\sum_{j\in\mathcal{B}_{r}}e^{V^{N}(c_{j})/\lambda_{r}}\Big)^{\lambda_{r}}}$
PCL $\sum_{r<s}\big(e^{V^{N}(c_{r})/\lambda_{rs}}+e^{V^{N}(c_{s})/\lambda_{rs}}\big)^{\lambda_{rs}}$ $\frac{\sum_{j\neq k}e^{V^{N}(c_{k})/\lambda_{kj}}\big(e^{V^{N}(c_{k})/\lambda_{kj}}+e^{V^{N}(c_{j})/\lambda_{kj}}\big)^{\lambda_{kj}-1}}{\sum_{r<s}\big(e^{V^{N}(c_{r})/\lambda_{rs}}+e^{V^{N}(c_{s})/\lambda_{rs}}\big)^{\lambda_{rs}}}$
GNL $\sum_{m=1}^{M}\Big(\sum_{j\in\mathcal{B}_{m}}(\alpha_{jm}e^{V^{N}(c_{j})})^{1/\lambda_{m}}\Big)^{\lambda_{m}}$ $\frac{(\alpha_{km}e^{V^{N}(c_{k})})^{1/\lambda_{m}}\Big(\sum_{j\in\mathcal{B}_{m}}(\alpha_{jm}e^{V^{N}(c_{j})})^{1/\lambda_{m}}\Big)^{\lambda_{m}-1}}{\sum_{r}\Big(\sum_{j\in\mathcal{B}_{r}}(\alpha_{jr}e^{V^{N}(c_{j})})^{1/\lambda_{r}}\Big)^{\lambda_{r}}}$
Let $G_{k}(\mathbf{y})=\frac{\partial G(\mathbf{y})}{\partial y_{k}}$, with $\mathbf{y}=(y_{1},\dots,y_{p})$. This derivative is crucial in determining how likely an individual is to choose alternative $c_{k}$ under the GEV model. If we evaluate it at $\mathbf{y}=\mathbf{e}^{\boldsymbol{\mu}^{N}}=\big(e^{V^{N}(c_{1})},\dots,e^{V^{N}(c_{p})}\big)$, the exponentiated systematic utilities, the term $G_{k}(\mathbf{e}^{\boldsymbol{\mu}^{N}})$ tells us how the attractiveness of option $c_{k}$ depends on the attractiveness of all the other available options. It adjusts for the fact that some choices are more closely related than others. For example, if two alternatives are close substitutes, a high utility for one will reduce the probability of choosing the other. Conversely, if options are weakly related, changing one has little effect on the probability of picking the other. Now, consider the following assumption: from the RUM (McFadden (1977)), we have:
$\pi_{gt}(c_{k})=\frac{\exp\Big(V^{N}_{gt}(c_{k})\Big)G_{k}(\mathbf{V^{N}_{gt}})}{G(\mathbf{V^{N}_{gt}})}=\frac{\exp\Big(\mu^{N}_{gt}(c_{k})\Big)G_{k}(\mathbf{\mu^{N}_{gt}})}{G(\mathbf{\mu^{N}_{gt}})}$
$q_{gt}^{N}(c_{k})=\pi_{gt}(c_{k})\cdot S^{N}_{gt}=\frac{\exp\Big(\mu^{N}_{gt}(c_{k})\Big)G_{k}(\mathbf{\mu^{N}_{gt}})}{G(\mathbf{\mu^{N}_{gt}})}\cdot S^{N}_{gt}.$
This leads to:
$\log q_{gt}^{N}(c_{k})$ $=$ $\mu_{gt}^{N}(c_{k})-\log\left(G(\mathbf{\mu^{N}_{gt}})\right)+\log(S_{gt}^{N})+\log(G_{k}(\mathbf{\mu^{N}_{gt}}))$
$\mathbb{E}[U_{g1}(c_{k})]-\mathbb{E}[U_{g0}(c_{k})]$ $=$ $\log q_{g1}^{N}(c_{k})-\log q_{g1}^{N}(c_{k})$
From there, we have that assumption 2 is equivalent to parallel trends in expected utilities. The result for utilities differences also follow.